OpenStudy (anonymous):

is x 1/2 the same as x squared?

OpenStudy (calculusxy):

Not necessarily. When you see that x is to the 1/2 power, you have to do multiply x 1/2 (0.5) times.

OpenStudy (anonymous):

x^(1/2) is the square root of x

OpenStudy (amonoconnor):

Umm, no. X^(1/2, or .5) = the square root of x. Think about it: to UNDO a square root, you square it... (x^(1/2))^2 = x^(2/2) = x^1 = x, so x to the one-half is the opposite, you could say of x squared.

OpenStudy (anonymous):

which, of course, is not the same as x^2 unless x = 1